1. In a box having 6 light switches, the probability of having a defective switch is 0.3 and the probability of having no defective switch is 0.6. Find the probability that there is at most 1 defective light switch in a box. 


1
Expert's answer
2021-05-10T13:09:06-0400

We have given that,

"n = 6"

"p = 0.3"

"q = 0.7"

We have to find the probability that there is at most 1 defective light switch in a box

Using the binomial distribution:


"P(X) = ^nC_r(p)^r(q)^{n-r}"


"P(X\le 1) = P(X=0)+P(X=1)"


"= ^6C_0(0.3)^0(0.7)^2+ ^6C_1(0.3)^1(0.7)^1"